OpenStudy (anonymous):

26,24,10 true or false

OpenStudy (anonymous):

I don't understand what this is asking. Is that the full question?

OpenStudy (anonymous):

Im trying to find if it is a right triangle

OpenStudy (anonymous):

Oh, well a right triangle is \[a^2+b^2=c^2\] So, \[26^2=10^2+24^2\] 676=100+576 676=676 The statement is true so it's correct.

OpenStudy (anonymous):

9,6,3 Is this a right triangle

OpenStudy (anonymous):

Same deal, \[9^2=3^2+6^2 \] 81=9+36 81=45 <----- Isn't a right triangle since the equation isn't true.

OpenStudy (anonymous):

10,8,6 Is this a right triangle

OpenStudy (anonymous):

I'm sure you can do this on your own now, right :$

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

The equation for a right triangle is \[a^2 +b^2=c^2\] "c" is always the biggest number, and "a" and "b" are always the other numbers.